# : Accelerating the Science of Language Models

### v3

@@ 2.1 OLMo Model and Architecture
2.2 Pretraining Data: Dolma
Despite progress in access to model parameters, pretraining datasets are still not as open. Pretraining data are often not released alongside open models (let alone closed models) and documentation about such data is often lacking in detail that would be needed to reproduce or fully understand the work. This has made it difficult to support certain threads of language model research, such as understanding how training data impacts model capabilities and limitations. To facilitate open research on language model pretraining, we built and released our pretraining dataset, Dolma—a diverse, multi-source corpus of 3T tokens across 5B documents acquired from 7 different data sources that are (1) commonly seen in large-scale language model pretraining and (2) accessible to the general public (Soldaini et al., 2024). Table 3 provides a high-level overview of the amount of data from each source.
Dolma is built using a pipeline of (1) language filtering, (2) quality filtering, (3) content filtering, (4) deduplication, (5) multi-source mixing, and (6) tokenization. We refer the reader to the Dolma report (Soldaini et al., 2024) for more details about its design principles, details about its construction, and a more detailed summary of its contents. The report provides additional analyses and experimental results from training language models on intermediate states of Dolma to share what we learned about important data curation practices, including the role of content or quality filters, deduplication, and mixing data from multiple sources. We keep documents from each source separate, both during curation as well as in the final release. We open-sourced our high-performance data curation tools; this toolkit can be used to further experiment on Dolma, reproduce our work, and enable fast and easy curation of pretraining corpora. Finally, we also open-sourced our WIMBD tool (Elazar et al., 2023) to help with dataset analysis.
Source | Doc Type |
web pages | 9,022 | 3,370 | 2,006 | ||||||||
code | 1,043 | 210 | 342 | ||||||||
web pages | 790 | 364 | 174 | ||||||||
social media | 339 | 377 | 80 | ||||||||
STEM papers | 268 | 38.8 | 57 | ||||||||
books | 20.4 | 0.056 | 5.2 | ||||||||
encyclopedic | 16.2 | 6.2 | 3.7 | ||||||||
Total | 11,519 | 4,367 | 2,668 |
2.3 Adaptation
Pretrained models are not always used as-is, but rather further fine-tuned to improve their performance, safety, and usability. Often models are first trained to follow instructions (Mishra et al., 2022; Wei et al., 2022; Sanh et al., 2022), and then further trained on human preferences (Ouyang et al., 2022) to improve the quality of their generations. We showcase the efficacy of using OLMo as a base model for further fine-tuning by training OLMo to be a general chat assistant following our Open Instruct (Tülu) data and training setup (Ivison et al., 2023). Our approach involves first performing instruction fine-tuning with a mixture of distilled and human-written instruction data and then further aligning the model with distilled preference data using Direct Preference Optimization (DPO) (Rafailov et al., 2023). We experimented with mixing the Tulu instruction data at the end of pretraining, as done in recent models such as DeepSeek-AI et al. (2024), but did not have conclusive findings.
2.4 Evaluation
We perform base model evaluation at two stages: online evaluation to make decisions for model design and offline evaluation to evaluate model checkpoints. For the offline stage, we use the Catwalk framework (Groeneveld et al., 2023), a publicly available evaluation tool with access to a wide range of datasets and task formats. Using Catwalk, we perform downstream evaluation as well as intrinsic language modeling evaluation on the new perplexity benchmark, Paloma (Magnusson et al., 2023).
In-Loop Training Ablations
Downstream Evaluation
Intrinsic Language Modeling Evaluation
To measure how OLMo-7B fits distributions of language beyond held-out training data, we use Paloma (Magnusson et al., 2023), a new perplexity benchmark that includes 585 different domains of text. Domains range from nytimes.com to r/depression on Reddit and are drawn from 18 separate data sources, such as C4 (Raffel et al., 2020), in stratified samples. This allows for more equal inclusion of text domains that are under-represented in their source corpora.
We aim not just to compare OLMo-7B against other models for best performance, but also to demonstrate how it enables fuller and more controlled scientific evaluations. OLMo-7B is the largest LM with explicit decontamination for perplexity evaluation. Following the approach described in Paloma, we remove any pretraining document with paragraphs leaked from Paloma evaluation data. Without decontamination, other models risk underestimating perplexity (i.e., overestimating the model’s out-of-sample fit). We also release intermediate checkpoints, allowing richer comparisons with two other models that release checkpoints, Pythia-6.9B (Biderman et al., 2023) and RPJ-INCITE-7B (Together Computer, 2023) (see Figure 2).
Adaptation Evaluation
We also follow our Open Instruct evaluation suite Wang et al. (2023); Ivison et al. (2023) to evaluate OLMo after instruction fine-tuning and DPO training using our We focus on evaluations around model chat capabilities and safety to showcase the efficacy of using OLMo as a base for further fine-tuning.
3 Training OLMo
This section describes our pretraining setup, including our distributed training framework (Section 3.1), optimizer settings (Section 3.2), data preparation (Section 3.3), and hardware (Section 3.4).
3.1 Distributed Training Framework
We train our models using the ZeRO optimizer strategy (Rajbhandari et al., 2019) via PyTorch’s FSDP framework (Zhao et al., 2023), which reduces memory consumption by sharding the model weights and their corresponding optimizer state across GPUs. At the 7B scale, this enables training with a micro-batch size of 4096 tokens per GPU on our hardware (see Section 3.4). For OLMo-1B and -7B models, we use a constant global batch size of approximately 4M tokens (2048 instances, each with a sequence length of 2048 tokens). For OLMo-65B model (currently training), we use a batch size warmup that starts at approximately 2M tokens (1024 instances), then doubles every 100B tokens until reaching approximately 16M tokens (8192 instances).
3.2 Optimizer
3.3 Data
Size | Peak LR | Betas | Epsilon | Weight Decay | Batch Size (tokens) |
1B | 4.0E-4 | (0.9, 0.95) | 1.0E-5 | 0.1 | 4M |
7B | 3.0E-4 | (0.9, 0.95) | 1.0E-5 | 0.1 | 4M |
65B* | 1.5E-4 | (0.9, 0.95) | 1.0E-5 | 0.1 | 2M 4M 8M 16M |
* At the time of writing our 65B model is still training.
OLMo-7B | LLaMA2-7B | OpenLM-7B | Falcon-7B | |
warmup steps | 5000 | 2000 | 2000 | 1000 |
peak LR | 3.0E-04 | 3.0E-04 | 3.0E-04 | 6.0E-04 |
minimum LR | 3.0E-05 | 3.0E-05 | 3.0E-05 | 1.2E-05 |
weight decay | 0.1 | 0.1 | 0.1 | 0.1 |
beta1 | 0.9 | 0.9 | 0.9 | 0.99 |
beta2 | 0.95 | 0.95 | 0.95 | 0.999 |
epsilon | 1.0E-05 | 1.0E-05 | 1.0E-05 | 1.0E-05 |
LR schedule | linear | cosine | cosine | cosine |
gradient clipping | global 1.0 | global 1.0 | global 1.0 | global 1.0 |
gradient reduce dtype | FP32 | FP32 | FP32 | BF16 |
@@ 3.3 Data
3.4 Hardware
4 Results
The checkpoint used for evaluating OLMo-7B is trained until 2.46T tokens on the Dolma (Soldaini et al., 2024) dataset with a linear learning rate decay schedule mentioned in Section 3.2. In our experiments, we find that tuning this checkpoint further on the Dolma dataset for 1000 steps with the learning rate linearly decayed to 0 boosts model performance on perplexity and end-task evaluation suites described in Section 2.4. We compare OLMo with other publicly available models including LLaMA-7B (Touvron et al., 2023a), LLaMA2-7B (Touvron et al., 2023b), MPT-7B (MosaicML NLP Team, 2023), Pythia-6.9B (Biderman et al., 2023), Falcon-7B (Almazrouei et al., 2023) and RPJ-INCITE-7B (Together Computer, 2023).
4.1 Downstream evaluation
Setup
Our core downstream evaluation suite (see Table 6) consists of: arc (both arc_easy and arc_challenge) (Clark et al., 2018), boolq (Clark et al., 2019), openbookqa (Mihaylov et al., 2018), sciq (Welbl et al., 2017), hellaswag (Zellers et al., 2019), piqa (Bisk et al., 2020), and winogrande (Sakaguchi et al., 2021). In Appendix A, we also report results on an additional set of auxiliary tasks outside of our core evaluation set that we found to have less stable performance trends (see Figure 4).
In all cases, we perform zero-shot evaluation using the rank classification approach popularized by Brown et al. (2020). Under this approach, candidate text completions (e.g., different multiple-choice options) are ranked by likelihood (usually normalized by some normalization factor), and prediction accuracy is reported. While Catwalk implements several common likelihood normalization strategies, including normalizing by number of tokens (per-token normalization) (Brown et al., 2020; Liang et al., 2022), by number of characters (per-character normalization) (Gao et al., 2023), as well as incorporating an answer’s unconditional likelihood (Brown et al., 2020), we selected the normalization strategies for each dataset separately. Specifically, we used unconditional normalization for arc and openbookqa, per-token normalization for hellaswag, piqa, and winogrande and no normalization for boolq, and sciq (i.e., tasks formulated as single token prediction tasks).
7B Models |
boolq |
piqa | sciq |
avg. | ||||||||||
Falcon | 47.5 | 70.4 | 74.6 | 75.9 | 53.0 | 78.5 | 93.9 | 68.9 | 70.3 | ||||||||||
LLaMA | 44.5 | 67.9 | 75.4 | 76.2 | 51.2 | 77.2 | 93.9 | 70.5 | 69.6 | ||||||||||
Llama 2 | 48.5 | 69.5 | 80.2 | 76.8 | 48.4 | 76.7 | 94.5 | 69.4 | 70.5 | ||||||||||
MPT | 46.5 | 70.5 | 74.2 | 77.6 | 48.6 | 77.3 | 93.7 | 69.9 | 69.8 | ||||||||||
Pythia | 44.1 | 61.9 | 61.1 | 63.8 | 45.0 | 75.1 | 91.1 | 62.0 | 63.0 | ||||||||||
RPJ-INCITE | 42.8 | 68.4 | 68.6 | 70.3 | 49.4 | 76.0 | 92.9 | 64.7 | 66.6 | ||||||||||
[] OLMo-7B | 48.5 | 65.4 | 73.4 | 76.4 | 50.4 | 78.4 | 93.8 | 67.9 | 69.3 |
Results
Table 6 summarizes the result of zero-shot evaluation of OLMo-7B and compares it against 6 other publicly available models of comparable size. We report results on 8 core tasks from our evaluation suite described in Section 2.4. On aggregate, OLMo-7B is competitive against all 6 publicly available model checkpoints in our comparison table.
In Figure 1 we plot the accuracy score progression of 8 core end-tasks. All tasks, except OBQA, show an upward trend in accuracy numbers as OLMo-7B is trained on more tokens. A sharp upward tick in accuracy of many tasks between the last and the second to last step shows us the benefit of linearly reducing the LR to 0 over the final 1000 training steps. See Table 9 in Appendix A for additional evaluation results and discussion.
4.2 Intrinsic language modeling evaluation
Setup
Results
The remaining subplots in Figure 2 provide more fine-grained analysis by reporting bits per byte separately for each of the 11 data sources that are combined in the aggregated Paloma metric. From this we see greater variation in sample efficiency, largely driven by the similarity of training and evaluation distributions. Notably, OLMo-7B fares well on evaluations predominated by Common Crawl, such as C4, though different ways of postprocessing Common Crawl are best fit by models trained with that specific data, such as Falcon-7B on Falcon RefinedWeb. Meanwhile, OLMo-7B is less sample efficient compared to other models on sources less related to scraped web text, such as WikiText-103, M2D2 S2ORC, and M2D2 Wikipedia. The RedPajama evaluation shows a similar pattern, perhaps as only 2 of its 7 domains are from Common Crawl, and Paloma weights domains within each source equally. Since heterogeneous data from curated sources like Wikipedia and ArXiv papers is much less abundant than scraped web text, maintaining sample efficiency for fit to these distributions of language will be challenging as pretraining corpora are scaled.
4.3 Adaptation Evaluation
Model MMLU AlpacaEval ToxiGen TruthfulQA 0-shot %win % Toxic %Info+True OLMo (base) 28.3 - 81.4 31.6 MPT Chat 33.8 46.8 0.1 42.7 Falcon Instruct 25.2 14.0 70.7 27.2 RPJ-INCITE Chat 27.0 38.0 46.4 53.0 Llama-2-Chat 46.8 87.3 0.0 26.3 Tülu 2 50.4 73.9 7.0 51.7 Tülu 2+DPO 50.7 85.1 0.5 - * [] OLMo +SFT 47.3 57.0 14.4 41.2 [] OLMo +SFT+DPO 46.2 69.3 1.7 52.0
Setup
We evaluate OLMo before adaptation, and after both the supervised fine-tuning and DPO training stage, focusing on the safety and chat evaluations used by Wang et al. (2023). We additionally compare to officially released instruction-tuned variants of the models from Table 6. We finally also compare to Tülu 2 models to compare against models trained using the same post-training data mixes and procedures.
Results
We find that instruction tuning considerably improves the performance and safety of OLMo, increasing MMLU performance by a wide margin and improving ToxiGen and TruthfulQA scores - especially after DPO training. Additionally, we find that OLMo outperforms most other chat variants after both initial instruction tuning (OLMo +SFT) and additional preference alignment (OLMo +SFT+DPO), highlighting both the strength of OLMo as a base model and the strength of the Tülu mix used to perform adaptation training. However, we find there is still a gap with Tülu 2, which is trained by applying the Tülu mix on Llama 2. This gap may be due to test set contamination in Llama 2777Touvron et al. (2023b) report that Llama 2 was pretrained on data contaminated with MMLU test data. and because the Tülu mix was primarily designed for Llama models - we will investigate the cause of this gap in future work. Overall, we see that OLMo greatly benefits from additional tuning and serves as a strong base model for downstream applications.
4.4 Power Consumption and Carbon Footprint
@@
Following previous literature (Strubell et al., 2019; Patterson et al., 2021; Wu et al., 2022; Dodge et al., 2022), we estimate the total energy consumed and carbon released while pretraining our models by calculating the total power consumption required for training, and then multiplying it by the carbon emission intensity of the power grid where the model was trained. While reporting these operational emissions is standard practice, it does not account for other sources of emissions such as the embodied emissions due to the manufacturing, transportation and disposal of hardware and datacenter infrastructure, lifetime operational emissions due to use, rebound effects, or other environmental impacts such as water consumption or mining. Thus our estimates should be viewed as lower bounds.
@@
We calculate the total power consumption for our models by measuring the power consumption of a single node every 25ms, calculating an average across the entire training run, and multiplying by the total number of nodes. We then account for the energy efficiency of the data center by multiplying the previous total by a power usage effectiveness (PUE) factor, which we set to 1.1, representing a conservative 10% energy consumption overhead typical of energy efficient datacenters.888https://www.nrel.gov/computational-science/measuring-efficiency-pue.html999https://www.google.com/about/datacenters/efficiency/ We estimate that pretraining our 7B models consumed 239 MWh of energy.
@@
To calculate carbon emissions, we multiply the total power consumption by a carbon intensity factor, measured in kg CO emitted per KWh, based on the physical location of the data center where each model was trained. The model trained on A100-40GB GPUs was trained in Australia, so we assume a carbon intensity factor of 0.610, the national average for Australia in 2022.101010https://www.cleanenergyregulator.gov.au/Infohub/Markets/Pages/qcmr/december-quarter-2022/Emissions-Reduction.aspx The model trained on MI250X GPUs was trained in the LUMI supercomputer, which runs on 100% renewable, carbon-neutral energy, so we assume a carbon intensity factor of 0. LUMI is powered entirely by hydroelectric power and some sources (Ubierna et al., 2022) measure the carbon intensity factor of hydroelectric power to be 0.024, which would imply total carbon emissions of 3.54 tCOeq.111111https://www.lumi-supercomputer.eu However, we rely on the official LUMI data for our calculations, and thus we estimate total pretraining emissions of 69.78 tCOeq.121212These metrics were in part collected using Carbonara’s AI agent and monitoring platform. Learn more at: https://trycarbonara.com In Table 12 we compare our models with other previously released models based on publicly available information.
@@
* LUMI runs entirely on hydroelectric power1212footnotemark: 12and some estimates (Ubierna et al., 2022) measure the intensity factor of hydroelectric power to be 0.024, implying total emissions of 3.54 tCOeq.
5 Artifacts Released
1.
The training and modeling code.131313https://github.com/allenai/OLMo
2.
The trained model weights for the 7B model,141414https://huggingface.co/allenai/OLMo-7B 7B-twin-2T,151515https://huggingface.co/allenai/OLMo-7B-Twin-2T and the 1B model.161616https://huggingface.co/allenai/OLMo-1B For all the models, we release not only the final model weights but also 500+ intermediate checkpoints at intervals of 1000 steps.
3.
Adapted OLMo-7B with instruction-tuning, 7B-SFT171717https://huggingface.co/allenai/OLMo-7B-SFT, and RLHF, 7B-Instruct181818https://huggingface.co/allenai/OLMo-7B-Instruct including its training and evaluation code and data using our Open Instruct191919https://github.com/allenai/open-instruct library (Wang et al., 2023; Ivison et al., 2023).
4.
The training data Dolma (Soldaini et al., 2024).202020https://huggingface.co/datasets/allenai/dolma
5.
Dolma’s toolkit to construct new datasets,212121https://github.com/allenai/dolma and WIMBD (Elazar et al., 2023) for dataset analysis.222222https://github.com/allenai/wimbd
6.
The evaluation code232323https://github.com/allenai/OLMo-Eval using Catwalk242424https://github.com/allenai/catwalk for downstream evaluation (Groeneveld et al., 2023) and Paloma252525https://paloma.allen.ai for perplexity-based evaluation (Magnusson et al., 2023).
7.
The complete set of metrics logged to Weights & Biases during training.262626https://wandb.ai/ai2-llm/OLMo-7B/reports/OLMo-7B--Vmlldzo2NzQyMzk5
We intend to follow up on this release with further training logs, ablations, and findings.
6 License
Our goal is to facilitate scientific development and empower the scientific community, so we favor permissive licenses that give users flexibility in using our resources and artifacts. As such, all code and weights are released under the Apache 2.0 License.272727http://www.apache.org/licenses/LICENSE-2.0 Some licenses used by other organizations for recent model releases prohibit using the outputs from their models to train artificial intelligence or machine learning systems, while we expressly allow users to do so. We also do not limit commercial use. We hope that our models can make other models better. We recognize that the risk for misuse of our models is relatively low since they are mainly designed as scientific artifacts not as products with broad public adoption (our models have not been adapted as chatbots). In addition, over the past year there have been a number of comparable models released with very permissive licenses, so using a more strict license for our work will not remove the overall risk in the field. We believe this tradeoff on the side of being more open is the best option.
7 Conclusion and Future Work
This technical report presents our first release of OLMo, a state-of-the-art, truly open language model and its framework to build and study the science of language modeling. Unlike most prior efforts that have only released model weights and inference code, we release OLMo and the whole framework, including training data and training and evaluation code. Soon, we will also release training logs, ablations, findings and Weights & Biases logs. We are also exploring the adaptation of OLMo with instruction tuning and different flavors of RLHF. We are going to release the adapted models as well as all of our model adaptation code and data.
We intend to continuously support and extend OLMo and its framework, and continue to push the boundaries of open LMs to empower the open research community. To that end, we look forward to bringing different model sizes, modalities, datasets, safety measures, and evaluations into the OLMo family. We hope this and future releases will empower and strengthen the open research community and inspire a new wave of innovation.
Author Contributions
OLMo would not have been possible without the help of our many teammates and collaborators. We list author contributions (in alphabetical order) below:
Contributors to pretraining dataset construction and tooling (Dolma) include Russell Authur, Iz Beltagy, Akshita Bhagia, Khyathi Chandu, Jesse Dodge, Yanai Elazar, Dirk Groeneveld, Rodney Kinney, Kyle Lo, Aakanksha Naik, Abhilasha Ravichander, Dustin Schwenk, Luca Soldaini, and Nishant Subramani.
Contributors to model training and architecture include Shane Arora, Iz Beltagy, Akshita Bhagia, Matthew E. Peters, Dirk Groeneveld, Ananya Harsh Jha, William Merrill, Jacob Morrison, Niklas Muennighoff, Dustin Schwenk, Saurabh Shah, Pete Walsh, and Mitchell Wortsman.
Contributors to evaluation suite and tooling include Akshita Bhagia, Arman Cohan, Pradeep Dasigi, Jesse Dodge, Dirk Groeneveld, Yuling Gu, Tushar Khot, Ian Magnusson, Kyle Richardson, Oyvind Tajford, and Pete Walsh.
Contributors to model adaptation include Iz Beltagy, Pradeep Dasigi, Jack Hessel, Hamish Ivison, Nathan Lambert, Valentina Pyatkin, Pete Walsh, and Yizhong Wang.
Contributors to license creation and risk assessment include David Atkinson, Jesse Dodge, Jennifer Dumas, Crystal Nam, and Will Smith.
The OLMo project was led by Hannaneh Hajishirzi and Noah A. Smith.
Acknowledgements
Appendix A Additional Evaluation
@@
In Figure 3 we provide results for each of the 7 data sources in Paloma [Magnusson et al., 2023] that are excluded from the combined metric in Figure 2. Some of these sources such as Pile [Gao et al., 2020] and ICE [Greenbaum and Nelson, 1996] are not publicly available at this time. Dolma 100 Programming Languages [Soldaini et al., 2024] consists of code data that is not supported by the decontamination approach used in Paloma. TwitterAAE [Blodgett et al., 2016], along with ICE, are datasets for targeted analyses of disparities in performance between different dialects and as such should be evaluated separately. And finally, the Manosphere, Gab, and 4chan corpora [Ribeiro et al., 2021, Zannettou et al., 2018, Papasavva et al., 2020] are intended to examine model fit to language from fringe online communities that are studied for prevalent hate speech and toxicity. Thus minimizing perplexity on these fringe corpora is not always desirable.
@@
Next, in Table 9, we provide results from zero-shot evaluation of OLMo-7B on 6 additional end-tasks apart from the 8 in our core evaluation suite. These tasks are headqa_en [Vilares and Gómez-Rodríguez, 2019], logiqa [Liu et al., 2020], mrpc [Dolan and Brockett, 2005], qnli [Wang et al., 2018], wic [Pilehvar and Camacho-Collados, 2018], and wnli [Wang et al., 2018].
headqa_en logiqa mrpc qnli wic wnli avg. Falcon-7B 38.6 23.7 62.8 49.8 49.5 47.9 45.4 LLaMA-7B 38.7 19.5 68.6 50.1 49.1 52.1 46.4 LLaMA2-7B 39.5 26.1 69.1 49.4 49.8 45.1 46.5 MPT-7B 37.4 22.9 67.7 52.1 48.1 47.9 46.0 Pythia-6.9B 40.1 21.5 65.4 53.8 55.0 38.0 45.6 RPJ-INCITE-7B 36.9 27.8 58.8 53.8 48.9 57.8 47.3 [] OLMo-7B 37.3 23.4 68.4 49.1 50.2 56.3 47.5
@@
Appendix B Adaptation Training Details

### v4

@@ Appendix A Training Settings
| Warmup steps | 5000 | 2000 | 2000 | 1000 | |
| Peak LR | 3.0E-04 | 3.0E-04 | 3.0E-04 | 6.0E-04 | |
| Minimum LR | 3.0E-05 | 3.0E-05 | 3.0E-05 | 1.2E-05 | |
| Weight decay | 0.1 | 0.1 | 0.1 | 0.1 | |
| Beta1 | 0.9 | 0.9 | 0.9 | 0.99 | |
| Beta2 | 0.95 | 0.95 | 0.95 | 0.999 | |
| Epsilon | 1.0E-05 | 1.0E-05 | 1.0E-05 | 1.0E-05 | |
| LR schedule | linear | cosine | cosine | cosine | |
| Gradient clipping | global 1.0 | global 1.0 | global 1.0 | global 1.0 | |
| Gradient reduce dtype | FP32 | FP32 | FP32 | BF16 | |
@@
Appendix B Power Consumption and Carbon Footprint
@@
Following previous literature (Strubell et al., 2019; Patterson et al., 2021; Wu et al., 2022; Dodge et al., 2022), we estimate the total energy consumed and carbon released while pretraining our models by calculating the total power consumption required for training, and then multiplying it by the carbon emission intensity of the power grid where the model was trained. While reporting these operational emissions is standard practice, it does not account for other sources of emissions such as the embodied emissions due to the manufacturing, transportation, and disposal of hardware and datacenter infrastructure, lifetime operational emissions due to use, rebound effects, or other environmental impacts such as water consumption or mining. Thus our estimates should be viewed as lower bounds.
@@
We calculate the total power consumption for our models by measuring the power consumption of a single node every 25ms, calculating an average across the entire training run, and multiplying by the total number of nodes. We then account for the energy efficiency of the data center by multiplying the previous total by a power usage effectiveness (PUE) factor, which we set to 1.1, representing a conservative 10% energy consumption overhead typical of energy efficient datacenters.101010https://www.nrel.gov/computational-science/measuring-efficiency-pue.html111111https://www.google.com/about/datacenters/efficiency/ We estimate that pretraining our 7B models consumed 239 MWh of energy.
@@
To calculate carbon emissions, we multiply the total power consumption by a carbon intensity factor, measured in kg CO2 emitted per KWh, based on the physical location of the data center where each model was trained. The model trained on A100-40GB GPUs was trained in Australia, so we assume a carbon intensity factor of 0.610, the national average for Australia in 2022.121212https://www.cleanenergyregulator.gov.au/Infohub/Markets/Pages/qcmr/december-quarter-2022/Emissions-Reduction.aspx The model trained on MI250X GPUs was trained in the LUMI supercomputer, which runs on 100% renewable, carbon-neutral energy, so we assume a carbon intensity factor of 0. LUMI is powered entirely by hydroelectric power and some sources (Ubierna et al., 2022) measure the carbon intensity factor of hydroelectric power to be 0.024, which would imply total carbon emissions of 3.54 tCO2eq.131313https://www.lumi-supercomputer.eu However, we rely on the official LUMI data for our calculations, and thus we estimate total pretraining emissions of 69.78 tCO2eq.141414These metrics were in part collected using Carbonara’s AI agent and monitoring platform. Learn more at: https://trycarbonara.com In Table 14 we compare our models with other previously released models based on publicly available information.
@@
* LUMI runs entirely on hydroelectric power1414footnotemark: 14and some estimates (Ubierna et al., 2022) measure the intensity factor of hydroelectric power to be 0.024, implying total emissions of 3.54 tCO2eq.
Appendix C Additional Evaluation
@@
In Figure 3 we provide results for each of the 7 data sources in Paloma (Magnusson et al., 2023) that are excluded from the combined metric in Figure 2. Some of these sources such as Pile (Gao et al., 2020) and ICE (Greenbaum and Nelson, 1996) are not publicly available at this time. Dolma 100 Programming Languages (Soldaini et al., 2024) consists of code data that is not supported by the decontamination approach used in Paloma. TwitterAAE (Blodgett et al., 2016), along with ICE, are datasets for targeted analyses of disparities in performance between different dialects and as such should be evaluated separately. And finally, the Manosphere, Gab, and 4chan corpora (Ribeiro et al., 2021; Zannettou et al., 2018; Papasavva et al., 2020) are intended to examine model fit to language from fringe online communities that are studied for prevalent hate speech and toxicity. Thus minimizing perplexity on these fringe corpora is not always desirable.
@@
Next, in Table 7, we provide results from zero-shot evaluation of OLMo-7B on 6 additional end-tasks apart from the 8 in our core evaluation suite. These tasks are headqa_en (Vilares and Gómez-Rodríguez, 2019), logiqa (Liu et al., 2020), mrpc (Dolan and Brockett, 2005), qnli (Wang et al., 2018), wic (Pilehvar and Camacho-Collados, 2018), and wnli (Wang et al., 2018).
| headqa_en | logiqa | mrpc | qnli | wic | wnli | avg. | |
| Falcon-7B | 38.6 | 23.7 | 62.8 | 49.8 | 49.5 | 47.9 | 45.4 |
| LLaMA-7B | 38.7 | 19.5 | 68.6 | 50.1 | 49.1 | 52.1 | 46.4 |
| LLaMA2-7B | 39.5 | 26.1 | 69.1 | 49.4 | 49.8 | 45.1 | 46.5 |
| MPT-7B | 37.4 | 22.9 | 67.7 | 52.1 | 48.1 | 47.9 | 46.0 |
| Pythia-6.9B | 40.1 | 21.5 | 65.4 | 53.8 | 55.0 | 38.0 | 45.6 |
| RPJ-INCITE-7B | 36.9 | 27.8 | 58.8 | 53.8 | 48.9 | 57.8 | 47.3 |
| [] OLMo-7B | 37.3 | 23.4 | 68.4 | 49.1 | 50.2 | 56.3 | 47.5 |
@@
Appendix D Adaptation Training Details